# Chirality Amplification and Deracemization in Liquid Crystals

### v1

@@ Acknowledgements.
Appendix A Supplementary Information
Dopant binding to a dense mesogen fluid
A simple model for the binding of chiral dopants to achiral mesogens may be provided by a Flory-Huggins type of fluid lattice model. The total number of lattice sites equals $M$. Let $\phi$ be the volume fraction of free mesogens, and $\phi_{*}$ and $\phi_{2}$ that of free dopant molecules and dimers of mesogens and dopants. The volume fractions add up to unity, $\phi+\phi_{*}+\phi_{2}=1$. Dimers take up two lattice sites, whilst free mesogens and free dopants each occupy only a single lattice site. The overall volume fraction of free and bound dopants is $\Phi_{*}$ and that of the free and bound mesogens is $\Phi=1-\Phi_{*}$. Hence, we have $\phi_{*}=\Phi_{*}-\frac{1}{2}\phi_{2}$ and $\phi=\Phi-\frac{1}{2}\phi_{2}$.
We ignore any net non-bonded interaction between mesogens, dopants and dimers, implying that the mesogen fluid acts as an ideal solvent to the dopants and dimers. We do account for a binding free energy $-g/k_{\mathrm{B}}T<0$ associated with each dimer. The Helmholtz free energy $F$ of the fluid mixtures in that case reads
@@ Appendix A Supplementary Information
where $k_{\mathrm{B}}T$ is the usual thermal energy, the first three terms account for the ideal entropy of mixing and the last for the binding of mesogens and dopants to form dimers. The optimal fraction dimers $\phi_{2}$ minimizes the free energy Eq. 7 ,
@@
It is useful to define the enantiomeric excess $\eta$, where we note that due to the presumed interconversion of P and M or $+$ and -- conformers, half the free mesogens are $+$ and half --. Let a chiral dopant bound to a mesogen turn this mesogen into a $+$ conformer. In that case $\phi_{+}=\frac{1}{2}\phi+\phi_{2}$ and $\phi_{-}=\frac{1}{2}\phi$, and $\eta=(\phi_{+}-\phi_{-})/\Phi=\phi_{2}/2\Phi$. Making use of the conservation of mass, we obtain from Eq. 8
@@
The free energy gain of dimerization is equal to $-\frac{1}{2}g\phi_{2}=-g\eta\Phi$ per lattice site. This means that the free energy gain per mesogen must be equal to $-g\hskip 2.84544pt\eta$. In the limit $r\ll 1$, we can replace $\eta$ by Eq. 11. In the same limit $r=\Phi_{*}/\Phi=\Phi_{*}/(1-\Phi_{*})\sim\Phi*$ and $\varepsilon=1/K\Phi\sim 1/K$. If the binding is strong and $-g\gg k_{\mathrm{B}}T$, we have $\varepsilon\ll 1$, and the binding free energy per mesogen becomes equal to $-g\Phi_{*}$ which can still be much smaller than the thermal energy if $\Phi_{*}$ is sufficiently small. This is the expression that we use in the main text, where we realize that for the lattice model, the volume fraction dopant is equal to its mole fraction, $X=\Phi_{*}$.
@@
According to our Maier-Saupe theory, the helical twisting power can be written as
$\beta=\frac{1}{1-\chi}\times\frac{q_{0}}{\pi}\times\left(\frac{g}{k_{\mathrm{B}}T}\right),$ (12)
@@
where the free energy gain $g=h-Ts$ can be written in terms of the enthalpy gain $h$ and the entropy loss $s$ of binding. For binding to occur we must heave $g>0$. If $h>0$ and $s>0$ then the binding is enthalpy-driven, if $h<0$ and $s<0$ the binding is entropy-driven. If all other factors do not depend on temperature, then $\beta$ decreases with increasing temperature if the binding is entropy driven.
@@
where $K=H_{K}-TS_{K}$ and $J=H_{J}-TS_{J}$ in terms of the enthalpy gains $H_{J,K}(T)$ and entropy losses $S_{J,K}(T)$ associated with the chiral and nematic interaction strengths $K\geq 0$ and $J\geq 0$. If the interaction strengths $J$ and $K$ are purely enthalpic in nature and $S_{J,K}=0$, as is usually presumed, then $q_{0}$ does not depend on the temperature. This is not what is usually seen experimentally albeit that the temperature dependence can be quite weak over a large temperature range, and for some chiral nematogens $q_{0}$ may in fact switch sign at some temperature [39]. To account for any temperature dependence, the Maier-Saupe theory for cholesterics has to be extended [13]. We argue, however, that since $J$ and $K$ represent interactions in a coarse-grained model in which microscopic degrees of freedom have been tacitly “integrated out”, these quantities should be viewed as free energies rather than energies (or enthalpies). Hence, depending on the sign and magnitude of $S_{J}$ and $S_{K}$, $q_{0}$ may increase or decrease with increasing temperature. This is similar in nature to how the Flory-Huggins interaction parameter of polymer solutions acquires its temperature dependence, and may increase or decrease with increase temperature depending on solvent and polymer type, leading to upper and lower critical solution temperatures.
@@
Figure S3: Cropped cross-section along simulation box diagonal $(+1,-1,+1)$ of a Monte Carlo simulation of size $N=480$ with $K/J=0.5$ and $T/T_{IN}=0.067$ after $2\times 10^{6}$ Monte Carlo steps showing the director configuration in the deracemized cholesteric phase. The colors red and blue indicate the left- and right-handed enantiomeric state of the model mesogens.
Figure S4: Chiral susceptibility in the racemic nematic phase, $\chi_{c}$, for $K/J=0.5$ and critical transition temperature, $T_{\mathrm{c}}/T_{IN}=0.134$. Drawn lines show the mean-field exponent of $\gamma=1$ and fitted exponent of $\gamma=1.036$.
@@
Fig. S4 shows a double-logarithmic plot of the chiral susceptibility, $\chi_{\eta}\sim|T-T_{\mathrm{c}}|^{-\gamma}$ as a function of the temperature $T-T_{\mathrm{c}}$ relative to the critical temperature $T_{\mathrm{c}}/T_{IN}=0.134$ in the racemic isotropic phase for the case $K/J=0.5$. The mean-field exponent of unity from fit of simulation data is shown in green and the fitted exponent is shown in blue.
@@
Fig. S6 shows final director and chirality configurations for completely annealed $N=128$ systems in various locations of the phase diagram. Fig. S6a) is in the racemic nematic phase and, while it shows small clusters of like chirality, there is no long-range chiral ordering and the nematic director is clearly visible. Fig. S6 b) shows a system that has just transitioned into a deracemized cholesteric phase with the chiral domains encompassing multiple pitches. Fig. S6 c) shows a system that hints at transitioning to the deracemized cholesteric blue phase as shown in Fig. S6 d). All director and chirality configurations were vizualised as the horizontal cross-section through the middle of the simulation volume on the $XZ$ plane with director field colored with the value of the Ising-like chiral variable.

### v2

@@ Acknowledgements.
Supplementary Information
Dopant interaction with mesogen enantiomers
A simple model for the interaction of chiral dopants to achiral mesogens may be provided by a Flory-Huggins type of fluid lattice model. The total number of lattice sites equals $M$. Let $\phi$ be the volume fraction of free mesogens, and $\phi_{*}$ and $\phi_{2}$ that of free dopant molecules in the fluid and dimeric complexes of mesogens and dopants, respectively. In reality, these dimers are not necessarily chemically (covalently) bound dimers but may in fact involve interaction complexes of a single dopant to more than a single nematogen. If so, then this merely renormalizes the binding free energy between dopant and mesogen to be introduced below. The free dopants interact much less strongly with the mesogens than those involved in what we refer to as the complexes. The volume fractions of all species add up to unity, $\phi+\phi_{*}+\phi_{2}=1$. Our model dimers take up two lattice sites, whilst free mesogens and free dopants each occupy only a single lattice site. The overall volume fraction of free and bound dopants is $\Phi_{*}$ and that of the free and bound mesogens is $\Phi=1-\Phi_{*}$. Hence, we have $\phi_{*}=\Phi_{*}-\frac{1}{2}\phi_{2}$ and $\phi=\Phi-\frac{1}{2}\phi_{2}$.
For the purpose of evaluating the net interaction free energy between strongly interacting mesogens and dopants, we may ignore any net attractive interaction between mesogens, dopants and dimers, and account only for their excluded volumes implying that the mesogen fluid acts as an ideal solvent to the dopants and the dimers. We do account for a binding free energy $g/k_{\mathrm{B}}T<0$ associated with each interaction complex. The Helmholtz free energy $F$ of the fluid mixtures in that case reads
@@
where $k_{\mathrm{B}}T$ is the usual thermal energy, the first three terms account for the ideal entropy of mixing and the last for the binding of mesogens and dopants to form dimers. Parenthetically, we not that if the dimers involve not one but $n$ mesogens, then the $1/2$ in the third term becomes $1/(n+1)$ and $g$ transforms to $ng$. This is renormalization we referred to above. The thermodynamics complexation does not change appreciably. Practically, it means $g$ is implied to have absorbed in it the unknown number $n$.
The optimal fraction dimers $\phi_{2}$ minimizes the free energy Eq. 7 ,
@@
It is useful to define the enantiomeric excess $\eta$, where we note that due to the presumed interconversion of P and M or ++ and - conformers, half the free mesogens are ++ and half -. Let a chiral dopant bound to a mesogen turn this mesogen into a ++ conformer. In that case $\phi_{+}=\frac{1}{2}\phi+\phi_{2}$ and $\phi_{-}=\frac{1}{2}\phi$, and $\eta=(\phi_{+}-\phi_{-})/\Phi=\phi_{2}/2\Phi$. Making use of the conservation of mass, we obtain from Eq. 8
@@
The free energy gain of complexation is equal to $\frac{1}{2}g\phi_{2}=-g\eta\Phi$ per lattice site. This means that the free energy gain per mesogen must be equal to $g\hskip 2.84544pt\eta$. In the limit $r\ll 1$, we can replace $\eta$ by Eq. 11. In the same limit $r=\Phi_{*}/\Phi=\Phi_{*}/(1-\Phi_{*})\sim\Phi*$ and $\varepsilon=1/K\Phi\sim 1/K$. If the binding is strong and $g\gg k_{\mathrm{B}}T$, we have $\varepsilon\ll 1$, and the binding free energy per mesogen becomes equal to $g\Phi_{*}$ which can still be much smaller than the thermal energy if $\Phi_{*}$ is sufficiently small. This is the expression that we use in the main text, where we realize that for the lattice model, the volume fraction dopant is equal to its mole fraction, $X=\Phi_{*}$.
In our model, we need not specify what physics underlies the binding free energy $g$, but it is nevertheless useful to spend a few words on it. In reality, the interaction process involves the alignment of one or more mesogens at the surface of the chiral dopant, such that it maximizes the van der Waals interaction with it. This arguably causes a loss of configurational entropy as well as van der Waals interaction with other mesogens as the “bound” mesogens may no longer be aligned with the bulk director field that in the model responds to the interaction with the dopant, as discussed in the main text. Part of the configurational entropy loss of the mesogens is that of the breaking of chiral symmetry. This entropy loss is dealt with explicitly in the full theory described in the main text. The dopant itself might also deform or have its conformational fluctuations affected by the interaction with the nematogens. This implies that $g$ has contributions from both the nematogens and the dopants.
@@
According to our Maier-Saupe theory, the dimensionless helical twisting power can be written as
$\beta=\frac{1}{1-\chi}\left(\frac{q_{0}g}{\pi k_{\mathrm{B}}T}\right),$ (12)
@@
where the free energy gain g=hTsg=h-Ts can be written in terms of the enthalpy gain $h$ and the entropy loss $s$ of binding. For binding to occur we must heave $g>0$. If $h>0$ and $s>0$ then the binding is enthalpy-driven, if $h<0$ and $s<0$ the binding is entropy-driven. If all other factors do not depend on temperature, then $\beta$ decreases with increasing temperature if the binding is entropy driven.
@@
where $K=H_{K}-TS_{K}$ and $J=H_{J}-TS_{J}$ in terms of the enthalpy gains $H_{J,K}(T)$ and entropy losses $S_{J,K}(T)$ associated with the chiral and nematic interaction strengths $K\geq 0$ and $J\geq 0$. If the interaction strengths $J$ and $K$ are purely enthalpic in nature and $S_{J,K}=0$, as is usually presumed, then $q_{0}$ does not depend on the temperature.
This is not what is usually seen experimentally albeit that the temperature dependence can be quite weak over a large temperature range, and for some chiral nematogens $q_{0}$ may in fact switch sign at some temperature [9]. To account for any temperature dependence, the Maier-Saupe theory for cholesterics has to be extended to include a higher order Legrendre polynomial in the interaction between the nematogens [46]. We argue, however, that since $J$ and $K$ represent interactions in a coarse-grained model in which microscopic degrees of freedom have been tacitly “integrated out”, these quantities should be viewed as free energies rather than energies (or enthalpies). Hence, depending on the sign and magnitude of $S_{J}$ and $S_{K}$, $q_{0}$ may increase or decrease with increasing temperature. This is similar in nature to how the Flory-Huggins interaction parameter of polymer solutions acquires its temperature dependence, and may increase or decrease with increase temperature depending on solvent and polymer type, leading to upper and lower critical solution temperatures.
@@
Figure S3: Cropped cross-section along simulation box diagonal (+1,1,+1)(+1,-1,+1) of a Monte Carlo simulation of size $N=480$ with $K/J=0.5$ and $T/T_{IN}=0.067$ after $2\times 10^{6}$ Monte Carlo steps showing the director configuration in the deracemized cholesteric phase. The colors red and blue indicate the left- and right-handed enantiomeric state of the model mesogens.
Figure S4: (Left) Chiral susceptibility in the racemic nematic phase, $\chi_{\eta}$, for $K/J=0.5$ and critical transition temperature, $T_{\mathrm{c}}/T_{IN}=0.134$. Drawn lines show the mean-field exponent of $\gamma=1$ and fitted exponent of $\gamma=1.036$. (Right) Chiral susceptibility in the racemic nematic phase, $\chi_{\eta}$, for $K/J=2.0$ and critical transition temperature, $T_{\mathrm{c}}/T_{IN}=1.2$. Drawn lines show the mean-field exponent of $\gamma=1$ and fitted exponent of $\gamma=0.933$.
@@
Figure S7: (a) Monte Carlo simulation results showing the director and chirality configuration of a $N=480$ simulation along the XZXZ axes at the midpoint of the $Y$-axis, with $K/J=0.5$ and $T/T_{IN}=0.067$ after $1\times 10^{5}$ Monte Carlo time steps. (b) Cropped inset showing a $100\times 100$ region of the $N=480$ simulation shown in (a) by the outlined region. (c) Cropped inset showing a $50\times 50$ region of the $N=480$ simulation shown in (b) by the outlined region.
@@
Fig. S4 shows two double-logarithmic plots of the chiral susceptibility, $\chi_{\eta}\sim|T-T_{\mathrm{c}}|^{-\gamma}$ as a function of the temperature $T-T_{\mathrm{c}}$ relative to the critical temperature. For the left plot, $K/J=0.5$, yielding critical temperature $T_{\mathrm{c}}/T_{IN}=0.134$ which is located in the racemic nematic phase. The mean-field exponent of unity from fit of simulation data is shown in green and the fitted exponent is shown in blue. The best fit gives a value of $\gamma=1.036\pm 0.029$. The right plot shows a similar fit above the critical endpoint, at $k/J=2.0$ which has a critical temperature $T_{\mathrm{c}}/T_{IN}=1.2$. The critical exponent $\gamma=0.933\pm 0.025$ is less than unity, suggesting that we cannot make a determination on whether the phase diagram contains a critical endpoint or a tricritical point, as $\gamma$ differs above and below this transition.
@@
Fig. S6 shows final director and chirality configurations for completely annealed $N=128$ systems in various locations of the phase diagram. Fig. S6 a) is in the racemic nematic phase and, while it shows small clusters of like chirality, there is no long-range chiral ordering and the nematic director is clearly visible. Fig. S6 b) shows a system that has just transitioned into a deracemized cholesteric phase with the chiral domains encompassing multiple pitches. Fig. S6 c) shows a system that hints at transitioning to the deracemized cholesteric blue phase as shown in Fig. S6 d). All director and chirality configurations were visualized as the horizontal cross-section through the middle of the simulation volume on the $XZ$ plane with director field colored with the value of the Ising-like chiral variable.
Fig. S7 shows the director and chirality configurations for a $N=480$ simulation. The interface between left- and right-handed cholesteric layers is sharp, even after only $1\times 10^{5}$ time steps, and is shown by insets focused on the layers far from the open boundaries. By examining the conditions away from the boundary, we can see that the open boundary conditions do not have a significant effect on the layer formation, nor the selection of the cholesteric pitch axis.
Supplementary Video
Video S1 shows an animation of the outer layer of chiral domains, as shown in Fig. S5 and Fig. 4c), to better show the coarsening behavior and the annihilation of layers of alternating chirality. This animation is of a singular simulation of size $480^{3}$ started in the racemic isotropic state and run at a constant temperature of $T/T_{IN}=0.067$ without annealing and a constant chiral interaction strength of $K/J=0.5$. Configurations of both the director and the Ising-like chiral variable were saved every $1\times 10^{5}$ Monte Carlo steps per spin for a total of $2\times 10^{6}$ steps and 20 frames of visualization.
Chiral susceptibility of 5CB
The nematogen 5CB has an isotropic-nematic transition temperature of $T_{\mathrm{IN}}=307$ K, and a nematic-to-crystal transition at $T_{\mathrm{Cr}}=296$[20]. This implies that the nematic is stable over a temperature range of $11$ K. Hence, for the chiral transition to be observed, $T_{\mathrm{c}}/T_{\mathrm{IN}}$ must be larger than $0.96$. Since it has not been observed, $T_{\mathrm{c}}/T_{\mathrm{IN}}$ must be smaller than this value. In Figure S8 we have plotted values of the chiral susceptibility $\chi_{\eta}$ as a function of the scaled temperature $T/T_{\mathrm{IN}}$ for a number of imagined values of $T_{\mathrm{c}}/T_{\mathrm{IN}}$ around the value of unity. Indicated with the dashed line is the crystal transition temperature. All curves diverge upon approach of the chiral transition temperature as $|T-T_{\mathrm{c}}|^{-1}$. Note that even if $T_{\mathrm{c}}<T_{\mathrm{IN}}$, the chiral transition still contributes to the chiral susceptibility and hence to the helical twisting power. For a discussion, we refer to the main text.
Figure S8: Dimensionless chiral susceptibility $\chi_{\eta}k_{\mathrm{B}}T_{\mathrm{IN}}$ as a function of the scaled temperature $T/T_{\mathrm{IN}}$, for different indicated values of the scaled chiral transition temperature $T_{\mathrm{c}}/T_{\mathrm{IN}}$. The dashed line indicates the crystal transition temperature of the compound 5CB. $T_{\mathrm{IN}}$ is the isotropic-nematic transition temperature, $T_{\mathrm{c}}$ the chiral transition temperature.
Helical twisting power: predictions compared
The prediction for the dimensionless helical twisting power $\beta$ from our Maier-Saupe theory is given in equation 12. Here, we seek comparison with the theory of Ferrarini and collaborators [15, 17, 16], which has been used to predict the helical twisting power of a class of chiral dopants from computer simulations [12]. The expression can be cast into the following form:
$\beta=\left(\frac{W\mathcal{C}}{2\pi K_{22}}\right),$
Here, $W$ can be seen interpreted as an anchoring surface energy of the director field of the nematic to the surface of the dopant, $\mathcal{C}$ a chirality order parameter describing the the coupling between the chiral surface of the molecule and its orientational ordering in the nematic host, and $K_{22}$ the twist elastic constant of the nematic host. The expression has been made non-dimensional by some unimportant microscopic length scale.
Arguably, $W$ and $\mathcal{C}$ depend on both the dopant and the nematogen, while $K_{22}$ is a property of the nematic host fluid. This suggests the following correspondence between both models: $W\mathcal{C}\leftrightarrow g$. The correspondence between $(1-\chi)^{-1}q_{0}$ and $1/K_{22}$ of Eqs. 12 and Helical twisting power: predictions compared is less obvious, as in the model of Ferrarini and collaborators the nematic is treated as a deformable background. However, since $q_{0}=K/2J$, and since $K_{22}\propto J$ [36], we do have $q_{0}\propto 1/K_{22}$. This is reassuring, and shows that both approaches have at least some common ground.
Clearly, both the factor $\chi\propto K^{2}$ as well as the $q_{0}\propto K$ involve the chiral interaction strength $K$ that is absent in the usual description of nematics. It is important to realize that the chiral shape theory is inherently macroscopic in nature, whilst our chiral fluctuation theory is microscopic in nature and does explicitly describe interaction of a director field with a chiral surface. It means that in the limit where $K\rightarrow 0$, our theory produces a zero helical twisting power. This, obviously, is drawback of our approach. Ideally, both theories should somehow be merged.
The simplest way to do that seems to be to replace $K_{22}$ in the macroscopic theory by a renormalized one that accounts for the chiral switching of the nematogens. Preliminary calculations show that within Maier-Saupe theory chiral interactions lower the twist constant according to $K_{22}\propto(1-\chi)$, and hence that the twist constant vanishes at the chiral transition temperature. This brings the two theories closer together. We intend to report on this in the near future [47].
Hysteresis
The Maier-Saupe model exhibits hysteresis for positive and negative values of the chiral biasing field $g$, provided the chiral interaction parameter $\chi$ is larger than unity, as shown in Figure S9. This is is indicative of a phase transition taking place for $\chi=1$. For $\chi>1$, both $\eta$ and $x$ change sign at zero field strength, $g=0$. The spinodals are state points at values of $g$ for which derivatives of these quantities diverge. The state points between the spinodals and the values at $g=0$ for which the slope is negative are thermodynamically unstable, and those for which the slopes are positive are metastable. The existence of metastable state points gives rise hysteresis phenomena similar to that found in ferromagnets.
Figure S9: Chiral order parameter $\eta$ (left) and scaled cholesteric wave number $x=q/q_{0}$ (right) as a function of the dimensionless field strength $g/k_{\mathrm{B}}T$. Lines indicate state points for which the free energy is stationary. Red: $\chi=2.0$ ; green: $\chi=1.5$; blue: $\chi=1.0$; orange: $\chi=0.5$; black: $\chi=0.1$